OpenStudy (anonymous):

help please!!!! The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). The probability that the urine sample contains the protein albumin is 0.58 (event B). The probability that a sample contains protein, given that albumin is detected, is 1. Which statement is true?

OpenStudy (anonymous):

Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) P(B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B) P(A). Events A and B are dependent because P(A|B) = P(A).

OpenStudy (anonymous):

\[P(A)=0.73\]\[P(B)=0.58\]\[P(A|B)=1\] Condition for independent events: \[P(AintersectionB)=P(A) \times P(B)\] and from conditional probability we have\[P(A|B)=\frac{P(AintersectionB)}{P(B)}\]\[\implies P(AintersectionB)=P(A|B) \times P(B)=1 \times 0.58=0.58\] Further we calculate \[P(A) \times P(B)=0.73 \times 0.58=0.4234\] Anything you can observe from these calculations?